Circular No. 8293

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004ab IN NGC 5054 L. A. G. Monard, Pretoria, South Africa, reports his discovery of a

supernova (mag 14.7) on unfiltered CCD images taken on Feb. 21.979 UT. The new object is located at α = 13^{h}16^{m}58^{s}.18, δ = −16^{o}37′52′′.7 (equinox 2000.0), which is 2′′ west and 11′′ north of the center of the galaxy NGC 5054. Nothing was visible at this location on images taken by Monard on 2003 Dec. 29.10 (limiting mag 18.5) and 2004 Feb. 1.15 (limiting mag 18.0) or on Digitized Sky Survey images. A confirming unfiltered CCD image taken by T. Vanmunster (Landen, Belgium) with a 0.35-m reflector on Feb. 24.13 shows SN 2004ab at mag 14.5.

T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that spectra of SN 2004ab, obtained by P. Berlind on Feb. 24.49 UT with the Mt. Hopkins 1.5-m telescope (range 370–750 nm) and by J. Huchra on Feb. 24.41 with the Cerro Tololo 4-m telescope (range 330–940 nm), show it to be a highly reddened type-Ia supernova about a week past maximum. Adopting the NASA/IPAC Extragalactic Database recession velocity of 1741 km/s for the host galaxy, the supernova expansion velocity is 10400 km/s for Si II (rest 635.5 nm). The spectra show a strong narrow Na I D absorption at the velocity of the host galaxy with an equivalent width of 0.33 nm. The dust maps of Schlegel et al. (1998, Ap.J. 500, 525) indicate a relatively low value for Galactic reddening of E(B − V) = 0.082. The spectral-feature age of the supernova (Riess et al. 1997, A.J. 114, 722) is 6 ± 2 days after maximum light.

S/2004 (17246) 1 P. M. Tamblyn, Southwest Research Institute (SwRI) and Binary As-

tronomy; W. J. Merline, C. R. Chapman, D. Nesvorný, and D. D. Durda, SwRI; C. Dumas, Jet Propulsion Laboratory; A. D. Storrs, Towson University; L. M. Close, University of Arizona; and F. Ménard, Observatoire de Grenoble, report the discovery on Jan. 14.9 UT, on six direct images (two sets of three images taken 20 min apart in time) made with the Hubble Space Telescope (+ ACS/HRC), of a satellite of minor planet (17246) 2000 GL_{74} (V ∼ 18.5). The satellite is clearly separated from the primary in five images. On Jan. 14.9195, the satellite was at separation 0′′.16 (projected separation 230 km) in p.a. 280^{o}. Using the average albedo of the Koronis family (∼ 0.21), to which (17246) belongs, the size of the primary is estimated to be 4.5 km. The brightness difference is ∼ 2 mag, giving an estimated diameter of the secondary of ∼ 2 km.

2004 February 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green